OpenStudy (anonymous):

Find f'(a) f(x)=3-2x+4x^2

OpenStudy (raden):

first, find f '(x). f(x) = 3-2x+4x^2, then f ' (x) = 0 - 2 + 8x = 8x - 2 therefore the value of f'(a) = ..... (just subtitute x = a in the equation of f'(x))